OpenStudy (anonymous):

Which of the following describes the graph of {x | x ≤ 1} U {x | x ≥ 4}? closed circles on 1 and 4 with a line segment between closed circles on 1 and 4 with arrows pointing outward the entire number line no solution i got D

OpenStudy (anonymous):

No, not D. \(\left\{x~|~x\le1\right\}\) and \(\left\{x~|~x\ge4\right\}\) is the same as \(1\le x\le4\).